the arithmetic mean of five numbers is 60. if the arithmetic mean of the first four numbers is 70, what is the fifth number?
Answer by mananth(16949)

arithmetic mean of five numbers = 60
so the sum of the numbers = 60*5 =300
..
first four numbers mean = 70
sum of these 4 numbers = 280.
so the 5th number is 300-280=20
the fifth number is 20